8. Let the random variable x represent the number of girls in a family with three...

8.

Let the random variable x represent the number of girls in a family with three children. Assume the probability of a child being a girl is 0.35 The table on the right describes the probability of having x number of girls. Determine whether the table describes a probability distribution. If it​ does, find the mean and standard deviation. Is it unusual for a family of three children to consist of three​ girls?

x   P(x)
0   0.275
1   0.444
2   0.239
3   0.042

Find the mean of the random variable. Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.

A. μ=

​(Round to two decimal places as​ needed.)

B. The table is not a probability distribution.

Determine whether the table describes a probability distribution. If it​ does, find the mean and standard deviation.

σ =

​(Round to two decimal places as​ needed.)

B. The table is not a probability distribution.

Is it unusual for a family of three children to consist of three​ girls?
Answer #1

Given probability distribition is,

X P(X) X * P(X) X2 * P(X)
0 0.275 0 0
1 0.444 0.444 0.444
2 0.239 0.478 0.956
3 0.042 0.126 0.378
X * P(X) 1.048 X2 * P(X) = 1.778

A) Mean of probability distribution of X is,

μ = E(X) 2 X * P(X) = 1.048 => μ = 1.05

B) standard deviation of X is,

2P(X) [E(X)]2-1.778 1.0482 0.8244

C) Yes, P(X=3) =0.042 < 0.05 therefore, the probability of having 3 girls is unusual.